ANALYSIS OF THE ν_{2} BAND OF THE FCO_{2} RADICAL: PRELIMINARY RESULTS

\underline{A. PERRIN}, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Universités Paris 12 et 7, 61 Av du Général de Gaulle, 94010 Créteil Cedex France; M. STŘIŽÍK, VŠB-Technical University of Ostrava, Faculty of Safety Engineering, Lumírova 13, CZ-70030, Ostrava 3 - Výškovice, Czech Republic, Institute of Thermomechanics, v.v.i., Academy of Sciences of the Czech Republic, Dolejškova 5, CZ-18200 Prague 8, Czech Republic; H. BECKERS, H. WILLNER, Inorg. Chemistry, University of Wuppertal, D-42119 Wuppertal, GERMANY; Z.ZELINGER, P.PRACNA, J. Heyrovský Institute of Physical Chemistry, Academy of Sciences of the Czech Republic, Dolejškova 3, CZ-18223 Prague 8, Czech Republic; V. NEVRLÝ, E. GRIGOROVÁ, VŠB-Technical University of Ostrava, Faculty of Safety Engineering, Lumírova 13, CZ-70030, Ostrava 3 - Výškovice, Czech Republic, Institute of Thermomechanics, v.v.i, Academy of Sciences of the Czech Republic, Dolejškova 5, CZ-18200 Prague 8, Czech Republic..

The unique fluoroformyloxyl radical (FCO_{2}) is assumed to participate in atmospheric processes such as the degradation of hydrofluorocarbons that have been considered as chlorofluorocarbon substitutes. Despite this atmospheric interest, the molecular and spectroscopic properties of FCO_{2} have not yet sufficiently been explored. The high resolution FT IR gas phase spectrum of the fluoroformyloxyl (FCO_{2}) radical was recorded in the 650 - 1500 cm^{−1} spectral range at the University of Wuppertal. Using this spectrum and the ground state parameters achieved recently^{a} we carried out the first high resolution study of the ν_{2} A-type band (C-F stretching mode) centered at 970.209 cm^{−1}. The analysis was difficult because the band is congested. In addition the spin doublets are difficult to identify except for high K_{a} values. However, we could take advantage of the fact that only K_{a}=odd values are observable for symmetry reasons. The line position calculation accounts for the spin rotation doubling and for the Fermi-type resonances linking the 2^{1} and 5^{2} spin rotation energy levels.

aKolesnikova, Varga, Beckers, Simeckova, Zelinger, Nova Striteska, Kania, Willner, and Urban, J. Chem Phys 128, 224 (2008)